# Efficient Semi-Supervised Adversarial Training via Latent Clustering-Based Data Reduction

### v2

@@ IV-A Strategic Unlabeled Data Selection
TABLE II: Runtime comparison of SSAT with synthetically generated unlabeled data across various settings on CIFAR-10.
Ratio $\bm{\alpha}$ Method Fine-Tuning (h) Generation (h) Selection (h) #Epochs Total Time (h) PGD (%)
@@
TABLE III: Performance of our DDPM fine-tuning strategies on CIFAR-10 under varying $\alpha$. The best robust accuracies are underlined.
Ratio $\bm{\alpha}$ Method Clean ($\%$) PGD ($\%$) AA ($\%$)
@@
IV-B Computational Benefits of Data Reduction
As suggested by the runtime comparisons demonstrated in Table I, utilizing smaller, carefully selected unlabeled data can significantly reduce the overall training time of SSAT while maintaining competitive robustness. We further visualize the learning curves of SSAT and discuss the computational benefits associated with our proposed methods.
Figure 1 depicts the learning curves of vanilla AT and SSAT algorithms on CIFAR-10 with varying ratios of unlabeled data selected from TinyImages. In our setup, one training epoch corresponds to processing 5050K data points, with $10$ PGD steps for training and $20$ for testing. When no additional data is utilized, the highest robust test accuracy is achieved after approximately $75$ epochs (Figure 1a). Incorporating $10\%$ additional data, whether through random selection or curated approaches, extends the optimal convergence point to approximately $100$ epochs (Figures 1b and 1c). In contrast, leveraging $100\%$ of the unlabeled data delays the achievement of peak accuracy to around $185$ epochs (Figure 1d). These findings suggest that including extra unlabeled data in SSAT increases computational demands. Our selection schemes reduce redundant information in the dataset, allowing SSAT to achieve high robustness performance with fewer training epochs.
Generally speaking, the size of the utilized unlabeled dataset is critical in determining the time required for SSAT to achieve peak robust accuracy. Larger unlabeled datasets usually extend the convergence times, whereas smaller datasets facilitate faster convergence, allowing the model to reach optimal accuracy earlier in training. However, continued training beyond this point often leads to robust overfitting. Early stopping offers a computational advantage in scenarios with limited datasets by preventing overfitting with reduced training durations. Nevertheless, early stopping proves less effective in SSAT when working with large-scale unlabeled datasets, as longer training is necessary to achieve peak performance. To mitigate these challenges, selecting a smaller yet strategically curated subset of data significantly reduces the dataset size, enabling faster convergence and minimizing overall training time. These findings underscore the computational benefits of our proposed data selection schemes, providing a practical solution to the challenges of SSAT with extensive datasets.
@@ IV-B Computational Benefits of Data Reduction
IV-C Guided DDPM Fine-Tuning
We study the performance of our guided DDPM fine-tuning methods as depicted in Algorithm 3 on CIFAR-10. Unlike our previous selection schemes, which required pre-generating the whole dataset, here we generate the same amount of unlabeled data matching the ratio $\alpha$. The data-mixing hyperparameters, namely $\alpha$ and $\beta$, are set consistently with those used in strategic selection. We employ a pretrained DDPM model trained on CIFAR-10 for these guided fine-tuning experiments. Table III presents the results across different $\alpha$ ratios, where our DDPM fine-tuning schemes achieve competitive clean accuracy and robustness performance, matching their selection counterpart. For instance, under the ratio scenario $\alpha=20\%$, leveraging LCG-KM to generate boundary-adjacent data achieves a PGD robust accuracy of $60.3\%$ when integrated in the SSAT framework, which matches the result of $60.2\%$ obtained by LCS-KM. Again, latent clustering with k-means consistently achieves the best robustness compared to the other alternatives. These findings mirror our findings from strategic selection, reinforcing that a small but informative generated set can yield performance comparable to a much larger dataset.
Moreover, we conduct a detailed runtime analysis across different data reduction schemes under the generative setup. In particular, we choose to present the best-performing methods with the optimal hyperparameters, namely LCS-KM and LCG-KM under the scenario of $\alpha=20\%$, and compare the end-to-end training times. We also report the baseline performance of SSAT, where the whole generated dataset ($\alpha=100\%$) and no generated data ($\alpha=0\%$) are used, respectively. The comparison results are summarized in Table II. For clarity, we separately report the computational time of DDPM fine-tuning, data generation, data selection, and the total SSAT training process. Our DDPM generator’s throughput is roughly $71$ images per second, so the approximate data generation times are 200200K images in $46$ minutes, 11M images in $3.9$ hours, and 100100M images in $16.2$ days. As a result, our guided DDPM finetuning (LCG-KM) approach reduces the total runtime from $19.1$ hours to $15.7$ hours when compared to LCS-KM. The improvement will be more pronounced when the total amount of data generated by DDPM is much larger. These results highlight the trade-off between robustness gains and computational cost. While generating millions of synthetic samples can enhance robustness, it is computationally expensive. Our guided fine-tuning approach offers a more resource-efficient alternative, reducing SSAT total runtime from $61.0$ hours to $15.7$ hours while preserving strong robustness.
@@ IV-C Guided DDPM Fine-Tuning
Figure 3: Visual comparison of selection techniques on TinyImages dataset in the latent space. Each subplot represents a different method: (a) PCS identifies points with the lowest classification confidence, highlighting areas where the model is most uncertain within the ten-class latent representation, (b) LCS-GMM illustrates probability contours from Gaussian Mixture Models, with selected points emphasizing regions of overlapping probabilities among the ten class clusters, and (c) LCS-KM highlights points selected near decision boundaries across ten classes based on k-means clustering in the latent space.

### v3

@@ V-A Strategic Unlabeled Data Selection
SVHN. Specifically for SVHN, we adopt a WRN-16-8 architecture. Employing our LCS-KM strategy with $10\%$ or $20\%$ of extra unlabeled data achieves PGD robust accuracies of $86.3\%$ and $86.6\%$, respectively, which matches and even slightly outperforms the $86.0\%$ robustness achieved using the entire unlabeled dataset. In contrast, random selection with $20\%$ of the data yields a robust accuracy of only $82.3\%$, underscoring the superior efficacy of informed selection strategies such as LCS-KM. The robust accuracies evaluated using AutoAttack demonstrate similar patterns compared to those using multi-step PGD Attacks. In addition, alternative methods, including PCS and LCS-GMM, consistently outperform random selection, confirming the effectiveness of our insights. Notably, the robustness of models trained using SSAT peaks when the boundary ratio parameter $\beta$ is inversely proportional to the selection ratio $\alpha$, suggesting that smaller $\beta$ values are favorable as the data inclusion proportion increases. We provide detailed sensitivity analyses of these hyperparameters in Section VI-C.
Our evaluations with 11M DDPM-generated data highlight that when using $10\%$ to $20\%$ of unlabeled data selected via LCS-KM, SSAT can achieve comparable robustness to the SSAT baseline. Since all SSAT algorithms listed under the column “DDPM-Generated” in Table I require the same amount of pre-generation time, we exclude it from the total runtime. In particular, training SSAT with reduced data requires only $75$ epochs to achieve the peak robust accuracy, while training with the full dataset requires up to $400$ epochs. Such a reduction in computation time underscores the practical advantages of the proposed data selection strategies, which are further elaborated in Section V-B. Besides, we observe in our experiments that using ground-truth labels reveals negligible differences compared to pseudo-labeled data, indicating that the primary performance gains stem from the selection strategy itself rather than label accuracy (see Table V in Appendix C).
CIFAR-10. We adopt a WRN-28-10 architecture for CIFAR-10 experiments. Across varying ratios $\alpha$ and data selection schemes, Table I reveals that selecting $20\%$ of the external data using LCS-KM attains PGD robust accuracy of $60.7\%$, close to the $62.5\%$ obtained with the entire dataset. In contrast, random selection with the same ratio achieves only $57.5\%$, underscoring the advantages of our strategic LCS technique. Similar to SVHN, we extend our analysis to CIFAR-10 with synthetically generated data. Using 11M DDPM-generated CIFAR-10-like images, we discover that using just $20\%$ of the generated data via LCS-KM results in performance nearly identical to that achieved with the full dataset. The above observations are also consistent across other settings, such as $\ell_{2}$ perturbations and a ResNet-18 model architecture (see Figure 9 and Table VI in the appendix for supporting evidence).
The above findings emphasize the adaptability of our data selection schemes, particularly for LCS-KM, under varying experimental conditions. When leveraging a small subset of selected external unlabeled data, optimal robust performance was achieved at $100$ epochs of training. In contrast, using the full external or generated datasets without selection requires extended training of $200$ and $400$ epochs, respectively, to reach peak performance. These results validate the efficacy of our latent clustering-based selection approach in enhancing semi-supervised adversarial training. The proposed method not only maintains the robust accuracy of the final model, $\theta_{\mathrm{final}}$, but also effectively accelerates the training process, underscoring its practical utility and scalability in SSAT frameworks.
V-B Computational Benefit of Data Reduction
As suggested by the runtime comparisons demonstrated in Table I, utilizing smaller, carefully selected unlabeled data can effectively reduce the overall training time of SSAT while maintaining competitive robustness. We further visualize the learning curves of SSAT and discuss the computational benefits associated with our proposed methods. Figure 1 depicts the learning curves of vanilla AT and SSAT algorithms on CIFAR-10 with varying ratios of unlabeled data selected from TinyImages. In our setup, one training epoch corresponds to processing 5050K data points, with $10$ PGD steps for training and $20$ for testing. When no additional unlabeled data is utilized, the highest test robust accuracy is achieved after approximately $75$ epochs (Figure 1a). Incorporating $10\%$ external unlabeled data, whether through random selection or curated approaches, extends the optimal convergence point to approximately $100$ epochs (Figures 1b and 1c). In contrast, leveraging $100\%$ of the unlabeled data delays the achievement of peak accuracy to around $185$ epochs (Figure 1d). These observations suggest that including extra unlabeled data in SSAT increases computational demands. Our selection schemes reduce redundant information in the dataset, allowing SSAT to achieve high robustness performance with fewer training epochs.
TABLE II: Performance of full SSAT ($\alpha=100\%$) with early stopping at epoch $100$ on CIFAR-10 under $\ell_{\infty}$ perturbations with $\epsilon=0.031$.
Data Source Schedule Clean ($\%)$ PGD ($\%)$ AA ($\%)$
External (500K) Cosine $84.1$ $54.0$ $50.4$
Stepwise $88.0$ $60.6$ $57.6$
DDPM (1M) Cosine $78.6$ $52.1$ $48.6$
Stepwise $83.0$ $58.5$ $55.9$
TABLE III: Runtime comparison of SSAT with synthetically generated unlabeled data across various settings on CIFAR-10.
$\bm{\alpha}$ Method Fine-Tuning (h) Generation (h) Selection (h) #Epochs Total Time (h) PGD (%)
@@ V-B Computational Benefit of Data Reduction
TABLE IV: Performance of our DDPM fine-tuning strategies on CIFAR-10 under varying $\alpha$. The best robust accuracies are underlined.
$\bm{\alpha}$ Method Clean ($\%$) PGD ($\%$) AA ($\%$)
@@
Generally, the size of the utilized unlabeled dataset is critical in determining the time required for SSAT to achieve peak robust accuracy. Larger unlabeled datasets usually extend the convergence times, whereas smaller datasets facilitate faster convergence, allowing the model to reach optimal accuracy earlier in training. Early stopping offers a computational advantage in scenarios with limited data by preventing overfitting with a reduced number of training epochs. Nevertheless, early stopping proves less effective in full SSAT when working with large-scale unlabeled datasets, as longer training is necessary to achieve peak performance. Table II summarizes the model performance on CIFAR-10 with respect to full SSAT when it is early stopped at the same training epoch as our methods. We tested two learning rate schedulers: cosine annealing and stepwise decay. The results show that the training process of full SSAT algorithms has not converged at the early-stopped training epoch. Importantly, even when comparing at the same number of training epochs, our LCS-KM method with $20\%$ DDPM-generated data achieves $60.2\%$ PGD robust accuracy at epoch $100$, outperforming early-stopped full SSAT by $1.7\%$. This demonstrates that our performance gain stems not only from reduced epochs but also from accelerated convergence enabled by strategically selecting more informative samples. Due to space limits, full results and detailed discussions are provided in Appendix D-C. These findings highlight the improved trade-offs of our methods in terms of computational costs, data efficiency, and robustness, offering a better solution to address the challenges of SSAT with extensive datasets.
@@
V-C Guided DDPM Fine-Tuning
We evaluate the performance of our guided DDPM fine-tuning methods on CIFAR-10. Unlike our selection-based schemes, which require pre-generating the entire dataset, here we generate the same amount of unlabeled data, matching the ratio $\alpha$. The data-mixing hyperparameters, namely $\alpha$ and $\beta$, are set consistently with those used in strategic selection. We employ a pretrained DDPM model trained on CIFAR-10 for these guided fine-tuning experiments. Table IV presents the results across different $\alpha$ ratios, where our DDPM fine-tuning schemes achieve competitive clean accuracy and robustness performance, matching their selection counterpart. For instance, under the ratio scenario $\alpha=20\%$, leveraging LCG-KM to generate boundary-adjacent data achieves a PGD robust accuracy of $60.3\%$ when integrated in the SSAT framework, which matches the result of $60.2\%$ obtained by LCS-KM. Again, latent clustering with k-means consistently achieves the best robustness compared to the other alternatives. These findings mirror our findings from strategic selection, reinforcing that a small but informative generated set of samples can yield performance comparable to a much larger dataset.
Moreover, we conduct a detailed runtime analysis across different data reduction schemes under the generative setup. In particular, we choose to present the best-performing methods with the optimal hyperparameters, namely LCS-KM and LCG-KM under the scenario of $\alpha=20\%$, and compare the end-to-end training times. We also report the baseline performance of SSAT, where the whole generated dataset ($\alpha=100\%$) and no generated data ($\alpha=0\%$) are used, respectively. The comparison results are summarized in Table III. For clarity, we separately report the computational time of DDPM fine-tuning, data generation, data selection, and the total SSAT training process. Our DDPM generator’s throughput is roughly $71$ images per second, so the approximate data generation times are 200200K images in $46$ minutes, 11M images in $3.9$ hours, and 100100M images in $16.2$ days. As a result, our guided DDPM finetuning (LCG-KM) approach reduces the total runtime from $19.1$ hours to $15.7$ hours when compared to LCS-KM. The improvement will be more pronounced when the total amount of data generated by DDPM is much larger. These results highlight the trade-off between robustness gains and computational cost. While generating millions of synthetic samples can enhance robustness, it is computationally expensive. Our guided fine-tuning approach offers a more resource-efficient alternative, reducing SSAT total runtime from $61.0$ hours to $15.7$ hours while preserving strong robustness.
@@ V-C Guided DDPM Fine-Tuning
Figure 3: Visual comparison of selection techniques on TinyImages dataset in the latent space. Each subplot represents a different method: (a) PCS identifies points with the lowest classification confidence, highlighting areas where the model is most uncertain within the ten-class latent representation, (b) LCS-GMM illustrates probability contours from Gaussian Mixture Models, with selected points emphasizing regions of overlapping probabilities among the $10$ class clusters, and (c) LCS-KM highlights points selected near decision boundaries across $10$ classes based on k-means clustering in the latent space.
VI Further Analyses
This section presents further analyses, including evaluations on a real-world medical dataset (Section VI-A), visualizations of selected unlabeled data (Section VI-B), and ablation studies on key hyperparameters (Section VI-C). Due to the page limit, additional experiments under different training schemes and perturbations, as well as the effect of the intermediate model in low-label regimes, are provided in Appendix D.
VI-A Application to Real-World Medical Data
We extend our study to a real-world healthcare application. COVIDGR [41] is used as the labeled dataset, and CoronaHack [16] serves as the source of the unlabeled data. Specifically, COVIDGR contains $852$ chest X-ray images, evenly distributed across two classes: $426$ COVID-19 positive (RT-PCR confirmed) and $426$ negative cases. It is then split into $80\%$ training and $20\%$ test sets. CoronaHack comprises $5910$ images with an imbalanced class distribution ($73\%$ positive, $27\%$ negative), split into an $89\%$ training set and an $11\%$ test set. We consider $\ell_{\infty}$ perturbations with $\epsilon=0.1$ and train models on a ResNet-18 architecture using SSAT by applying our LCS-KM strategy to select unlabeled CoronaHack data with varying ratios (see Appendix 3 for more details).
Figure 2 demonstrates the results, which again confirm the effectiveness of our selection scheme. For instance, using LCS-KM to select $10\%$ of the unlabeled data achieves a robust accuracy of $56\%$, compared to $53\%$ when the same proportion is chosen randomly (Figures 2a and 2b). In comparison, using a smaller unlabeled dataset leads to faster convergence, underscoring the benefits of early stopping. Specifically, the best performance of $56\%$ is reached at epoch $36$ when using $10\%$ of the data selected with LCS-KM (Figure 2b). Figure 2c shows that incorporating $20\%$ of the data with LCS-KM improved the accuracy to $57\%$ at epoch $47$. In contrast, training on the entire unlabeled dataset requires $60$ epochs to achieve a peak robust accuracy of $58\%$ (Figure 2d). These findings confirm that using a smaller, but strategically selected, unlabeled data subset can achieve robustness comparable to that of the entire dataset. These results highlight the practical relevance of our data reduction scheme, particularly for applications where labeled data is limited yet robust model performance is essential.
(a) CIFAR-10 ($\alpha=20\%$)
(b) SVHN ($\alpha=10\%$)
(c) Fine-Tuning Epochs $S$
(d) Regularization Strength $\lambda$
Figure 4: Figures (a) and (b): Illustration of SSAT performance using LCS-KM by varying boundary ratio parameters $\beta$ with respect to DDPM-generated CIFAR-10 data with $\alpha=20\%$ and 531531K extra SVHN data with $\alpha=10\%$. Figures (c) and (d): Clean and robust accuracy curves of SSAT with LCG-KM on CIFAR-10 with $\alpha=10\%$ by varying: (a) the number of fine-tuning epochs $S$, and (b) the regularization strength parameter $\lambda$.
VI-B Comparative Visualization
To facilitate a deeper understanding, we perform a comparative visualization of the selected unlabeled data on CIFAR-10. We first extract the latent representations from the penultimate layer of a pre-trained WRN-28-10 intermediate model. For visualization, we employ a two-step dimension reduction that balances the retention of global structure and the visualization of local relationships. In particular, we first perform Principal Component Analysis (PCA) [14]. By reducing the dimensionality of the latent features to $50$, PCA retains the majority of the variance within the data. Then, we apply t-Distributed Stochastic Neighbor Embedding (t-SNE) [46] to project the $50$-dimensional PCA features onto a $2$-dimensional space. Unlike PCA, which emphasizes preserving variance, t-SNE mainly aims to maintain the local relationships among data points. Finally, we visualize the set of unlabeled data selected by our methods under $\alpha=10\%$. Figure 3 shows the visualization results, where each color represents the label class predicted by the intermediate model for the corresponding unlabeled point. The highlighted black dots indicate the data points selected by the respective selection strategy.
Specifically, PCS adopts a purely confidence-based selection strategy, choosing the points with the globally lowest confidence without considering their distribution within or across classes. Although this approach is appealing for its simplicity, it can inadvertently prioritize noisy outliers. As the number of classes grows and data complexity increases, this scattershot selection tends to yield a more disorganized set of points (Figure 3a). In comparison, LCS-GMM and LCS-KM yield more structured sampling patterns. While LCS-GMM can roughly capture regions where classes overlap, it assumes that each cluster follows approximately Gaussian distributions, an assumption that may fail in real-world settings. As a result, LCS-GMM’s contours become misaligned with the true underlying structure, leading to less accurate localization of essential boundary points (Figure 3b). A closer look at the selected samples also reveals that the corresponding region is broader and less concentrated on the decision boundary, making it harder to capture truly critical samples for SSAT. Our best-performing LCS-KM method emphasizes the importance of leveraging the latent-space clustering to identify boundary-adjacent data points in a more accurate and balanced manner. By partitioning the latent embeddings into multiple k-means clusters, LCS-KM is able to locate more points with low classification confidence within each cluster (Figure 3c). Compared with LCS-GMM, a notable pattern is that the unlabeled data picked by LCS-KM are aligned more linearly. This coincides with the piecewise linear characteristic of neural networks’ decision boundary, especially for the last layer. We hypothesize that this local linear property may explain why our latent clustering-based approach with k-means consistently achieves the best performance compared to the other two alternatives.
VI-C Hyperparameter Sensitivity
We conduct ablations to study the sensitivity of the hyperparameters involved in our methods The detailed setup and the ablations on the unlabeled data ratio $\alpha$ and the per-batch ratio of extra to original data $\gamma$ are deferred to Appendix C-A.
Ratio of Boundary Data $\bm{\beta}$. Concentrating exclusively on data points near the decision boundary can result in an excessive number of points close to the boundary and too few farther away, leading to overfitting. To mitigate this, we choose a combination of points, with some near the boundary and others further away, where the ratio is determined by $\beta$. Figure 4a illustrates how robust accuracy varies with $\beta$ on CIFAR-10 using $20\%$ generated data from DDPM. In this scenario, the optimal robust accuracy is achieved when $\beta$ is $0.4$, meaning $40\%$ of the selected points are near the boundary, while $60\%$ are farther away. Figure 4b illustrates how robust accuracy varies with $\beta$ on SVHN using $10\%$ extra data from the same dataset. Due to a smaller amount of external data, selecting points near the boundary results in a lower risk of overfitting. As a result, the highest robust accuracy is observed when $\beta$ is $0.6$ with $60\%$ of the points near the boundary and $40\%$ farther away. Therefore, the ideal value of $\beta$ depends on the amount of external data added and the ratio of external to original data. When more external data is added, using a smaller $\beta$ typically yields better robust accuracy. Conversely, choosing a larger $\beta$ is favorable when there is less extra unlabeled data.
Guided DDPM Fine-Tuning ($S$ & $\lambda$). We conduct ablations on our best-performing LCG-KM approach by varying the number of fine-tuning epochs $S$ and the regularization strength parameter $\lambda$ on CIFAR-10 under $\alpha=10\%$. Figure 4c illustrates our ablation results on $S$. We vary the number of fine-tuning epochs $S\in\{5,15,25,35,45\}$ with $\lambda=0.5$, and report the final clean and robust accuracies after the whole SSAT process. We observe that choosing a small number of epochs for fine-tuning DDPM (e.g., 10102020 epochs) is usually sufficient to achieve the best robustness performance, likely because the starting DDPM is already pre-trained on CIFAR-10. Training for too many epochs can lead to a degradation in image quality and loss of diversity (see Figure 5 in the appendix for visualizations), thereby adversely influencing the final robust accuracies. Figure 4d shows the results for the regularization strength parameter $\lambda\in\{0.1,0.3,0.5,0.7\}$ with fixed epoch number $S=15$. Our findings indicate that selecting $\lambda$ within the range of $[0.4,0.6]$ yields the best performance, striking a balance between the quality of the generated images and their proximity to the model’s decision boundary.
VII Conclusion and Future Work
In this work, we developed various data reduction schemes to improve the efficiency of SSAT. We illustrated that strategically selecting or generating more boundary-adjacent data points can significantly improve the data efficiency of SSAT while preserving its robustness advantages. In particular, our latent-based k-means clustering approach performs the best, reducing both computational costs and memory requirements.
Limitation & Future Direction. While LCS-KM consistently achieves the best performance, its success requires careful hyperparameter tuning; future work can leverage AutoML algorithms to automate this process. In addition, theoretical analysis of how to solve the optimization problems (Equations 4 and 6) in a more principled way would be helpful. Finally, extending our guided DDPM fine-tuning with training-free sampling-based guidance [11] or advanced generative models [38, 13] is a promising direction. Further discussions are provided in Appendix D-D. Despite these challenges, we believe our work can serve as an important initial step towards more efficient and scalable development of robust models.
Appendix A Algorithm Pseudocode
Algorithms 1-3 present the pseudocode of our data reduction schemes, corresponding to prediction confidence-based selection (PCS), latent clustering-based selection (LCS), and guided DDPM-based fine-tuning, as introduced in Section IV.
Appendix B Detailed Experimental Setup
SVHN. SVHN is naturally divided into a core training set comprising approximately 7373K images and an extra training set of around 531531K images. Initially, the model is trained on the 7373K labeled images. To evaluate the generalizability of our selection schemes, we apply them to a synthetic dataset generated using a DDPM model [17].
Configuration. For model training, we adopt a WideResNet 16-8 (WRN-16-8) architecture [48] in all our SVHN experiments. We generate adversarial examples using PGD attacks exactly as implemented by Zhang et al. [50], with step size $0.007$, $10$-step PGD attack iterations, and $\ell_{\infty}$ perturbation magnitude $\epsilon=0.015$. Hyperparameters are set the same as in Carmon et al. [5] except for the number of epochs. For SVHN, we use a training batch size of $128$. All of our experiments that utilize all of the extra 531K data are run for $200$ epochs, and the experiments with 11M generated data are run for $400$ epochs. To prevent overfitting and reduce computational complexity, we run our experiments with selected data for $75$ epochs with early stopping. Following TRADES [50], we employ an SGD optimizer with a weight decay factor of $5\cdot 10^{4}$ and an initial learning rate of $0.1$, along with a stepwise learning rate scheduler with a decaying factor of $10$ at epochs $75$ and $90$.
Algorithm 1 Prediction Confidence-Based Selection (PCS)
4:  Initialize $\mathcal{A}_{u}$ as empty set
6:  Sort $\mathcal{S}_{u}$ by ascending order of $\mathrm{Conf}(\bm{x})$
Algorithm 2 Latent Clustering-Based Selection (LCS)
4:  Initialize $\mathcal{A}_{u}$ as empty set
6:  if using LCS-KM then
10:  else if using LCS-GMM then
11:   Fit a GMM with $k$ components to $\{\bm{z}\}_{\bm{x}\in\mathcal{S}_{u}}$
12:   Compute GMM posterior probability $\bm{p}$ for each $\bm{z}$
15:  end if
Algorithm 3 Guided DDPM Fine-Tuning
3:  Initialize $\mathcal{G}_{u}$ as empty set
5:  Initialize $\theta$ as $\theta_{\mathrm{pre}}$
6:  for $s=1,2,\ldots,S$ do
10:   if using PCG then
12:   else if using LCG-KM then
14:   else if using LCG-GMM then
16:   end if
18:  end for
Evaluation. For the attack evaluation to calculate the robust accuracy using PGD attack, we keep the parameters similar to those of Carmon et al. [5] for better comparison. We use a step size $0.005$, and a number of attack steps of $K=100$. We evaluate models at $\epsilon=0.015$, which is the same value used during training. We also use Auto Attack with the same $\epsilon$ value as in the PGD attack.
CIFAR-10. The CIFAR-10 dataset has 50K labeled images. We use the $80$ Million Tiny Images (80M-TI) dataset, of which CIFAR-10 is a manually labeled subset, to obtain extra data. However, most images in 8080M-TI do not correspond to CIFAR-10 image categories. Carmon et al. [5] used an 11-way classifier to distinguish CIFAR-10 classes and an 11th “non-CIFAR-10” class using a WideResNet 28-10 model. Each class selected an additional 5050K images from 80M-TI using the model’s predicted scores to create a 500500K pseudo-labeled dataset, which we used in our experiments. We train the intermediate model using 5050K labeled data and utilize this model to select data from either the 500500K pseudo-labeled data or a synthetic dataset. We perform the same experiments as in SVHN, except that we do not have the ground-truth labels of the additional unlabeled dataset. We also conducted further experiments using 11M synthetic data generated by the DDPM model [17]. For our fine-tuning experiments, we start with a publicly available pretrained DDPM model [20] and further fine-tune it for a few epochs to generate the desired images, which are subsequently used for our final training.
Configuration. For our main CIFAR-10 experiments, we use a WideResNet 28-10 (WRN-28-10) [48] architecture to train models. Similar to SVHN experiments, we set step size $0.007$, PGD attack iterations as $10$, and $\ell_{\infty}$ perturbation magnitude $\epsilon=0.031$. For our adversarial training with $\ell_{2}$ perturbation, we use $\epsilon=128/255$. In section D-A, when $\epsilon=4/255$, we use $0.0035$ as the PGD step size. For $\epsilon=16/255$, we use $0.015$ as the PGD step size. Hyperparameters used are the same as in Carmon et al. [5] except for the number of epochs. Here, we use a training batch size of $256$. The experiments using all the 500500K pseudo-labeled data are run for $200$ epochs, and using 1M generated data are run for $400$ epochs. For our experiments with limited data, we get the best results at $100$ epochs, where we early stop the training process. We employ an SGD optimizer with a weight decay factor of $5\cdot 10^{4}$ and an initial learning rate of $0.1$, along with a stepwise scheduler with a decaying factor of $10$ at epochs $75$ and $90$, respectively.
Evaluation. The PGD attack evaluation is conducted similarly to that of Carmon et al. [5] for fair comparisons. We use a step size $\alpha=0.01$, and a number of attack steps $K=40$. We consider $\ell_{\infty}$ perturbations with $\epsilon=0.031$, the same as the perturbation magnitude we used during training. For the models trained with $\ell_{2}$ perturbation, we evaluate with $\ell_{2}$ perturbation magnitude $\epsilon=128/255$. We also use Auto Attack with the same $\epsilon$ value as in the PGD attack.
Medical Application. For experiments on the medical dataset outlined in Section VI-A, we train models using a ResNet-18 architecture [19], and consider $\ell_{\infty}$ perturbations with $\epsilon=0.1$. We set the attack step size to $0.02$, and the number of PGD steps to $10$. The final model is trained using SSAT with the Adam optimizer [25] and a learning rate of $0.001$. We evaluate model robustness with PGD attacks with a step size of $0.02$ and $20$ attack steps.
Figure 5: Visualization of generated images with various fine-tuning epochs $S$.
Appendix C Additional Experiments